# BNSR-Invariants of Surface Houghton Groups

### v1

@@ 3 The Stein–Farley complex $X_{n}$ is CAT(0)
First, observe that the equality $[L_{g},\varphi\circ\rho_{i}]=[L_{g},\varphi^{\prime}\circ\rho_{j}]$ means that $h=\rho_{i}^{-1}\varphi^{-1}\varphi^{\prime}\rho_{j}$ takes $L_{g}$ to itself, and is rigid elsewhere; also, the equality $[L_{g+1},\varphi]=[L_{g+1},\varphi^{\prime}]$ means that $\varphi^{-1}\varphi^{\prime}(L_{g+1})=L_{g+1}$. Assume $j\neq n$, and consider the piece $B^{1}_{j}$; let $\alpha$ be the boundary curve it shares with $\mathcal{O}$. What does $h$ do to $\alpha$? First, $\rho_{j}$ deforms it into the center and stretches it into the $n$th end, making it an essential curve in $L_{g}$. Then $\varphi^{-1}\varphi^{\prime}$ moves it around to some essential curve in $L_{g+1}$. Finally, $\rho_{i}^{-1}$ pushes and stretches it towards the $i$th end. As $B_{j}^{1}$ is not in $L_{g}$, it must be sent to a piece, and $h(\alpha)$ must be the intersection of this piece with $L_{g}$. As all representatives here do not permute ends, it must be that $h(\alpha)=\alpha$. If $i\neq j$, then $\rho_{i}$ fixes $\alpha$, so $\rho_{i}(\alpha)\neq\varphi^{-1}\varphi^{\prime}\rho_{j}(\alpha)$, and $h(\alpha)\neq\alpha$. Hence we must have that $i=j$. In the case where $j=n$, equality is obvious. ∎
Now, we seek the following equality: $[L_{g+2},\varphi\circ\rho_{k}]=[L_{g+2},\varphi^{\prime}\circ\rho_{k}]$. That is, we wish to show that $\rho_{k}^{-1}\varphi^{-1}\varphi^{\prime}\rho_{k}$ takes $L_{g+2}$ to itself, and is rigid elsewhere. We already know that this holds for $\rho_{i}^{-1}\varphi^{-1}\varphi^{\prime}\rho_{i}$. The rigidity here occurs precisely when $\varphi^{-1}\varphi^{\prime}$ does nothing untoward to whatever single piece gets pushed into the $n$th end, so it doesn’t matter which index we choose! (If we have chosen the index $n$, there is no pushing. To resolve this, simply choose a different height $g+2$ vertex to work with. They can’t share ends, as they are in squares with lower vertices, so we can apply Lemma 3.5.)
One small detail remains: we have a bottom vertex connecting to one of the height $g+1$ vertices, but does it connect to the other height $g+1$ vertices? Suppose not: then the cube between $[L_{g+3},\varphi]$ and $[L_{g},\varphi\circ\rho_{i}\rho_{j}\rho_{k}]$ contains a square differing from one of the original 3 squares only at its height $g+1$ vertex (that is, at its lowest vertex). This however means that we would have two squares agreeing on 3 vertices, disagreeing on the 4th, which would contain a collapsed square. As this cannot occur (Lemma 3.6), our height $g$ vertex connects to each of the height $g+1$ vertices we started with.
{tikzpicture}
@@ 4 $\Sigma$-invariants of $P\mathcal{H}_{n}$
The methods of this section are taken from [Zar16] and [Zar20]. They need only minimal modification to work in this setting, and are presented partly for the sake of having the entire argument in one place. We begin by noting that the abelianization of $P\mathcal{H}_{n}$ is $\mathbb{Z}^{n-1}$, and the abelianization map is $(\chi_{1},\dots,\chi_{n-1})$ (see Section 6 of[ABKL23]). As $\chi_{1}+\cdots+\chi_{n}=0$, any character $\chi$ can be written (up to renumbering the ends of $\Sigma_{n}$) in ascending standard form, i.e. $\chi=a_{1}\chi_{1}+\cdots+a_{n}\chi_{n}$, with $a_{1}\leq\cdots\leq a_{m(\chi)}<a_{m(\chi)+1}=\cdots=a_{n}=0$. We shall henceforth assume all characters are written in such form. Observe that $m(\chi)$ is defined to be the maximal index so that $a_{m(\chi)}<a_{n}=0$. Then our goal in this section is to prove the following.
@@
Let $\chi$ be a non-zero character of $P\mathcal{H}_{n}$. Then $[\chi]\in\Sigma^{m(\chi)-1}(P\mathcal{H}_{n})\setminus\Sigma^{m(\chi)}(P\mathcal{H}_{n})$.
@@ 4.1 Inclusion
For any non-zero character $\chi\in\operatorname{Hom}(P\mathcal{H}_{n},\mathbb{R})$, we have $[\chi]\in\Sigma^{m(\chi)-1}(P\mathcal{H}_{n})$.
@@
Let $h=(\chi,f):Y\rightarrow\mathbb{R}\times\mathbb{R}$ be a Morse function. If $Y$ is $(m-1)$-connected and for every vertex $v\in Y^{\chi<q}$ the ascending link $\operatorname{lk}_{Y}^{h\uparrow}(v)$ is $(m-1)$-connected, then $Y^{q\leq\chi}$ is $(m-1)$-connected.
From now on, we write $X$ for $X_{n}$. The role of $Y$ shall be played by sublevel subccomplexes $X^{f\leq q}$ for sufficiently large $q$. Let $\chi=a_{1}\chi_{1}+\cdots+a_{n}\chi_{n}$ be a non-trivial character of $P\mathcal{H}_{n}$, written in ascending standard form. The function $h=(\chi,f):Y\rightarrow\mathbb{R}^{2}$ is a Morse function. Observe that between any two adjacent vertices, each basis character differs by $1$, 00, or $-1$. We examine $h$-ascending links in $Y$.
@@
Let $v=[L_{g},\varphi]$ be a vertex in $Y$. An adjacent vertex $w=[L_{g\pm 1},\psi]$ is in the $h$-ascending link of $v$ if and only if one of the two following conditions hold:
@@
We shall need the following fact: Let $v\in X_{n}$ be a vertex. If $f(v)\geq 2n$, then the descending link of $v$ is $(n-2)$-connected.
Now set $q=3n-2$, so $Y=X^{f\leq 3n-2}$. We note that this index is one greater than that used for the regular Houghton group. This is because the bounds obtained for connectivity of descending links obtained in [Lee12] and [ABKL23] differ by one (they are $12n-$ 1 and $2n$ respectively, see [Lee12, Theorem 3.52] and [ABKL23, Section 5.2]). We have the following:
@@
We have that $f(v)$ is between 00 and $23n-$ 2. Suppose that $f(v)\leq 2n+m(\chi)-2$. Writing $v=[L_{g},\varphi]$, we have that there are $n-m(\chi)$ indices for which $v^{\prime}=[L_{g+1},\varphi\circ\rho_{i}^{-1}]$ is $h$-ascending. As $(2n+m(\chi)-2)+(n-m(\chi))=3n-2$, the entire $f$-ascending link of $v$ in $X$ is contained in $Y$. Thus, the $f$-ascending part of the $h$-ascending link is an $(n-m(\chi)-1)$-simplex, which is contractible, so that $\operatorname{lk}_{Y}^{h\uparrow}(v)$ is contractible.
Suppose instead that $22n+m(\chi)-1\leq f(v)\leq 3n-$ 2, and thus that $Y$ does not contain the entire $f$-ascending part of the $h$-link of $v$. We still have its $(3n-f(v)-3)$-skeleton, which, being a skeleton of an $(n-m(\chi)-1)$-simplex, is $(3n-f(v)-4)$-connected. As $f(v)\geq 2n$, we have that the entire $f$-descending part of the $h$-ascending link in is $Y$. This is isomorphic to the $f$-descending link of a vertex of the same $f$-height as $v$ in $X_{m(\chi)}$, which we know is $(m(\chi)-2)$-connected so long as $f(v)\geq 2m(\chi)$. As $f(v)\geq 2n+m(\chi)-1$, we need that $2n-1\geq m(\chi)$, which is certainly true. The join is now $((3n-f(v)-3)+(m(\chi)-1))$-connected. As $f(v)\leq 3n-2$, we have that $(3n-f(v)+m(\chi)-4)\geq m(\chi)-2$, and thus that $\operatorname{lk}_{Y}^{h\uparrow}(v)$ is $(m(\chi)-2)$-connected. ∎
@@ 4.2 Exclusion
For any non-zero character $\chi\in\operatorname{Hom}(P\mathcal{H}_{n},\mathbb{R})$, we have $[\chi]\notin\Sigma^{m(\chi)}(\mathcal{H}_{n})$.
@@
Let $X$ be a CW-complex covered by subcomplexes $(X_{i})_{i\in I}$ and let $L$ be the nerve of the cover. Let $n\geq 1$. Suppose that any non-empty intersection $X_{i_{1}}\cap\cdots\cap X_{i_{r}}$ is $(n-r)$-connected. Then $H_{k}(X)\cong H_{k}(L)$ for all $k\leq n-1$, and $H_{n}(X)$ surjects onto $H_{n}(L)$.
@@
Recall that a subcomplex $Z$ of a CAT(0) cube complex $Y$ is locally combinatorially convex if every link in $Z$ of a vertex $z$ is a full subcomplex of the link of $z$ in $Y$, and combinatorially convex if it is connected and locally combinatorially convex. It is known that combinatorially convex implies CAT(0), hence contractible. In particular, this applies to connected components of locally combinatorially convex subcomplexes.
@@
It suffices to show that each $X^{\chi_{i}\leq 0}$ is locally combinatorially convex. Given a pair of adjecent vertices, we can write them as $v=[L_{g},\varphi]$ and $w=[L_{g+1},\varphi\circ\rho_{j}^{-1}]$ for some $j$. Then $\chi_{i}(w)-\chi_{i}(v)=\delta_{i,j}$. Thus, if $C$ is a cube containing $v$, and $w_{1},\dots,w_{k}$ are the vertices of $C$ adjacent to $v$, then the maximum and minimum values of $\chi_{i}$ on $C$ lie in $\{\chi_{i}(v),\chi_{i}(w_{1}),\dots,\chi_{i}(w_{k})\}$. Thus, whenever $v\in X^{\chi_{i}\leq 0}$ and all these $w_{j}$’s lie in the link of $v$ in $X^{\chi_{i}\leq 0}$, then the cube $C$ lies in $X^{\chi_{i}\leq 0}$. This implies that the link of $v$ in $X^{\chi_{i}\leq 0}$ is a full subcomplex of the link of $v$ in $X$. ∎
As in [Zar20], we have the immediate corollary that intersections of blankets are blankets for the union of the sets of indices. What follows is essentially identical to Zaremsky’s approach for the Houghton groups, reproduced here with minor changes for completeness’ sake. Recall now the more general statement of the Morse lemma, Lemma 2.11. For notational convenience, we write $X_{f\leq k}$ for $X^{f\leq k}$, and $X_{f\leq k}^{t\leq\chi}$ for the intersection $X_{f\leq k}\cap X^{t\leq\chi}$.
@@
If $X_{f\leq 3n-2}^{0\leq\chi}$ is not $(m(\chi)-1)$-connected, then $[\chi]\in\Sigma^{m(\chi)}(P\mathcal{H}_{n})^{c}$.
@@
If $[\chi]\in\Sigma^{m(\chi)}(P\mathcal{H}_{n})$, then the filtration $(X_{f\leq 3n-2}^{t\leq\chi})_{t\in\mathbb{R}}$ is essentially $(m(\chi)-1)$-connected. Every $h$-ascending link of a vertex in $X_{f\leq 3n-2}$ is $(m(\chi)-2)$-connected, so for any $s\leq t$ the inclusion $X_{f\leq 3n-2}^{t\leq\chi}\hookrightarrow X_{f\leq 3n-2}^{s\leq\chi}$ induces an isomorphism in $\pi_{k}$ for $k\leq m(\chi)-2$, and a surjection in $\pi_{m(\chi)-1}$. By assumption, for any $t$ there is some $s\leq t$ such that this inclusion induces a trivial map in $\pi_{k}$ for $k\leq m(\chi)-1$, implying that $X_{f\leq 3n-2}^{s\leq\chi}$ is $(m(\chi)-1)$-connected. Rescaling if necessary, we can assume that $s\in\chi(P\mathcal{H}_{n})$, and thus we can translate to obtain $X_{f\leq 3n-2}^{s\leq\chi}\cong X_{f\leq 3n-2}^{0\leq\chi}$, so that $X_{f\leq 3n-2}^{0\leq\chi}$ is itself $(m(\chi)-1)$-connected. ∎
@@
To be non-empty, such an intersection can include at most one term $Y_{i}^{\alpha}$ for each $i$, and can thus be written $Y=Y_{i_{1}}^{\alpha_{1}}\cap\cdots\cap Y_{i_{r}}^{\alpha_{r}}$, with the $i_{j}$ all pairwise distinct. Let $Z=Z_{i_{1}}^{\alpha_{1}}\cap\cdots\cap Z_{i_{r}}^{\alpha_{r}}$, so that $Y=Z\cap X_{f\leq 3n-2}^{0\leq\chi}$. We apply Morse theoretic techinques to $Z$, this time using Lemma 2.11. As $Z$ is an intersection of blankets, it is a blanket, and thus contractible. Given adjacent vertices $w=[L_{g},\varphi]$ and $v=[L_{g+1},\varphi\circ\rho_{i}^{-1}]$ with $v\in Z$, we have that $w\in Z$. Thus, for any vertex of $Z$, the entire $f$-descending link is in $Z$. As this is $(n-2)$-connected for $f(v)\geq 2n$, we see that $Z_{f\leq 3n-2}$ is $(n-2)$-connected, and so is certainly $(m(\chi)-2)$-connected. Now consider $Y$ as $Z_{f\leq 3n-2}^{0\leq\chi}$. As before, the $h$-ascending link of a vertex $v$ is a join between its $f$-ascending and $f$-descending parts. The latter is in $Z$ for the same reasons as above; the former is in $Z$ because it consists of directions $i$ where $m(\chi)+1\leq i\leq n$, on which the considered $\chi_{i_{j}}$ are constant. Thus, the $h$-ascending link of $v$ is in $Z_{f\leq 3n-2}$. As $Z_{f\leq 3n-2}$ is $(m(\chi)-2)$-connected, the Morse lemma tells us that $Y$ is $(m(\chi)-2)$-connected. ∎
Let $L$ be the nerve of the covering of $X_{f\leq 3n-2}^{0\leq\chi}$ by the $Y_{i}^{\alpha}$. Since $[\chi]\in\Sigma^{m(\chi)-1}(P\mathcal{H}_{n})$, we know that $X_{f\leq 3n-2}^{0\leq\chi}$ is $(m(\chi)-2)$-connected, so by the Strong Nerve Lemma $L$ is $(m(\chi)-2)$-connected. The final missing piece is to prove that $L$ is not $(m(\chi)-1)$-acyclic.
@@
It remains to show that $Y_{i}^{1}\neq Y_{i}^{2}$ for each $i$. It suffices to show that $Z_{i}^{1}\neq Z_{i}^{2}$. If these are equal (call them $Z_{i}$), then we can connect $v_{0}$ to $v_{i}$ via a path in $Z_{i}$. In $X$, one could assume that such a path is one on which $f$ first strictly increase, then strictly decreases; since $Z_{i}$ is combinatorially convex, this property holds for $Z_{i}$ as well. Since the path lies in $Z_{i}$, $\chi_{i}$ is non-positive on the whole path. By the uniqueness of ascending edges, this is a path of the form
{tikzcd}

### v2

@@ 3 The Stein–Farley complex $X_{n}$ is CAT(0)
First, observe that the equality $L_{g},\varphi\circ\rho_{i}]=[L_{g},\varphi^{\prime}\circ\rho_{j}]$ means that $h=\rho_{i}^{-1}\varphi^{-1}\varphi^{\prime}\rho_{j}$ takes $L_{g}$ to itself, and is rigid elsewhere; also, the equality $L_{g+1},\varphi]=[L_{g+1},\varphi^{\prime}]$ means that $\varphi^{-1}\varphi^{\prime}(L_{g+1})=L_{g+1}$. Assume $j\neq n$, and consider the piece $B^{1}_{j}$; let $\alpha$ be the boundary curve it shares with $\mathcal{O}$. What does $h$ do to $\alpha$? First, $\rho_{j}$ deforms it into the center and stretches it into the $n$th end, making it an essential curve in $L_{g}$. Then $\varphi^{-1}\varphi^{\prime}$ moves it around to some essential curve in $L_{g+1}$. Finally, $\rho_{i}^{-1}$ pushes and stretches it towards the $i$th end. As $B_{j}^{1}$ is not in $L_{g}$, it must be sent to a piece, and $h(\alpha)$ must be the intersection of this piece with $L_{g}$. As all representatives here do not permute ends, it must be that $h(\alpha)=\alpha$. If $i\neq j$, then $\rho_{i}$ fixes $\alpha$, so $\rho_{i}(\alpha)\neq\varphi^{-1}\varphi^{\prime}\rho_{j}(\alpha)$, and $h(\alpha)\neq\alpha$. Hence we must have that $i=j$. In the case where $j=n$, equality is obvious. ∎
Now, we seek the following equality: $L_{g+2},\varphi\circ\rho_{k}]=[L_{g+2},\varphi^{\prime}\circ\rho_{k}]$. That is, we wish to show that $\rho_{k}^{-1}\varphi^{-1}\varphi^{\prime}\rho_{k}$ takes $L_{g+2}$ to itself, and is rigid elsewhere. We already know that this holds for $\rho_{i}^{-1}\varphi^{-1}\varphi^{\prime}\rho_{i}$. The rigidity here occurs precisely when $\varphi^{-1}\varphi^{\prime}$ does nothing untoward to whatever single piece gets pushed into the $n$th end, so it doesn’t matter which index we choose! (If we have chosen the index $n$, there is no pushing. To resolve this, simply choose a different height $g+2$ vertex to work with. They can’t share ends, as they are in squares with lower vertices, so we can apply Lemma 3.5.)
One small detail remains: we have a bottom vertex connecting to one of the height $g+1$ vertices, but does it connect to the other height $g+1$ vertices? Suppose not: then the cube between $L_{g+3},\varphi]$ and $L_{g},\varphi\circ\rho_{i}\rho_{j}\rho_{k}]$ contains a square differing from one of the original 3 squares only at its height $g+1$ vertex (that is, at its lowest vertex). This however means that we would have two squares agreeing on 3 vertices, disagreeing on the 4th, which would contain a collapsed square. As this cannot occur (Lemma 3.6), our height $g$ vertex connects to each of the height $g+1$ vertices we started with.
$L_{g+3},\varphi][$$L_{g+3},\varphi^{\prime}][$$L_{g+3},\varphi^{\prime\prime}][$$L_{g+1},\varphi\circ\rho_{i}\rho_{j}][$$L_{g+1},\varphi^{\prime}\circ\rho_{i}\rho_{k}][$$L_{g+1},\varphi^{\prime\prime}\circ\rho_{j}\rho_{k}][$$L_{g+2},\varphi\circ\rho_{i}][$$L_{g+2},\varphi^{\prime}\circ\rho_{i}][$$L_{g+2},\varphi\circ\rho_{j}][$$L_{g+2},\varphi^{\prime\prime}\circ\rho_{j}][$$L_{g+2},\varphi^{\prime}\circ\rho_{k}][$$L_{g+2},\varphi^{\prime\prime}\circ\rho_{k}][$$L_{g+3},\varphi][$$L_{g+3},\varphi^{\prime}][$$L_{g+3},\varphi^{\prime\prime}]$
@@ 4 $\Sigma$-invariants of $P\mathcal{H}_{n}$
The methods of this section are taken from [Zar16] and [Zar20]. They need only minimal modification to work in this setting, and are presented partly for the sake of having the entire argument in one place. We begin by noting that the abelianization of $P\mathcal{H}_{n}$ is $\mathbb{Z}^{n-1}$, and the abelianization map is $(\chi_{1},\dots,\chi_{n-1})$ (see Section 6 of [ABKL23]). As $\chi_{1}+\cdots+\chi_{n}=0$, any (non-trivial) character $\chi$ can be written (up to renumbering the ends of $\Sigma_{n}$) in ascending standard form, i.e. $\chi=a_{1}\chi_{1}+\cdots+a_{n}\chi_{n}$, with $a_{1}\leq\cdots\leq a_{m(\chi)}<a_{m(\chi)+1}=\cdots=a_{n}=0$. We shall henceforth assume all characters are written in such form. Observe that $m(\chi)$ is defined to be the maximal index so that $a_{m(\chi)}<a_{n}=0$. Then our goal in this section is to prove the following.
@@
Let $\chi$ be a non-zero character of $P\mathcal{H}_{n}$. Then $\chi]\in\Sigma^{m(\chi)-1}(P\mathcal{H}_{n})\setminus\Sigma^{m(\chi)}(P\mathcal{H}_{n})$.
@@ 4.1 Inclusion
For any non-zero character $\chi\in\operatorname{Hom}(P\mathcal{H}_{n},\mathbb{R})$, we have $\chi]\in\Sigma^{m(\chi)-1}(P\mathcal{H}_{n})$.
@@
Let $h=(\chi,f):Y\rightarrow\mathbb{R}\times\mathbb{R}$ be a Morse function. If $Y$ is (m1)(m-1)-connected and for every vertex $v\in Y^{\chi<q}$ the ascending link $\operatorname{lk}_{Y}^{h\uparrow}(v)$ is (m1)(m-1)-connected, then $Y^{q\leq\chi}$ is (m1)(m-1)-connected.
From now on, we write $X$ for $X_{n}$. The role of $Y$ shall be played by sublevel subcomplexes $X^{f\leq q}$ for sufficiently large $q$. Let $\chi=a_{1}\chi_{1}+\cdots+a_{n}\chi_{n}$ be a non-trivial character of $P\mathcal{H}_{n}$, written in ascending standard form. The function $h=(\chi,f):Y\rightarrow\mathbb{R}^{2}$ is a Morse function. Observe that between any two adjacent vertices, each basis character differs by $1$, 0, or 1-1. We examine $h$-ascending links in $Y$.
@@
Let $v=[L_{g},\varphi]$ be a vertex in $Y$. An adjacent vertex $w=[L_{g\pm 1},\psi]$ is in the $h$-ascending link of $v$ if and only if one of the two following conditions holds:
@@
We shall need the following fact (see [ABKL23, Section 5.2]): Let $v\in X_{n}$ be a vertex. If $f(v)\geq 2n$, then the $f$-descending link of $v$ is (n2)(n-2)-connected. Note that the version of this statement for the regular Houghton groups uses $f(v)\geq 2n-1$ (see [Lee12, Theorem 3.52]); this discrepancy will result in most of the numbers in the following proposition being either one above or one below the corresponding numbers in [Zar16].
Now set q=3n2q=3n-2, so $Y=X^{f\leq 3n-2}$. We have the following:
@@
We have that $f(v)$ is between 0 and 3n23n-2. Suppose that $f(v)\leq 2n+m(\chi)-2$. Writing $v=[L_{g},\varphi]$, we have that there are $n-m(\chi)$ indices $i$ for which $v^{\prime}=[L_{g+1},\varphi\circ\rho_{i}^{-1}]$ is $h$-ascending. As $(2n+m(\chi)-2)+(n-m(\chi))=3n-2$, the entire $f$-ascending link of $v$ in $X$ is contained in $Y$. Thus, the $f$-ascending part of the $h$-ascending link is an $(n-m(\chi)-1)$-simplex, which is contractible, so that $\operatorname{lk}_{Y}^{h\uparrow}(v)$ is contractible.
Suppose instead that $2n+m(\chi)-1\leq f(v)\leq 3n-2$, and thus that $Y$ does not contain the entire $f$-ascending part of the $h$-link of $v$. We still have its (3nf(v)3)(3n-f(v)-3)-skeleton, which, being a skeleton of an $(n-m(\chi)-1)$-simplex, is (3nf(v)4)(3n-f(v)-4)-connected. As $f(v)\geq 2n$, we have that the entire $f$-descending part of the $h$-ascending link in is $Y$. This is isomorphic to the $f$-descending link of a vertex of the same $f$-height as $v$ in $X_{m(\chi)}$, which we know is $(m(\chi)-2)$-connected so long as $f(v)\geq 2m(\chi)$. As $f(v)\geq 2n+m(\chi)-1$, we need that $2n-1\geq m(\chi)$, which is certainly true. The join is now $((3n-f(v)-3)+(m(\chi)-1))$-connected. As $f(v)\leq 3n-2$, we have that $(3n-f(v)+m(\chi)-4)\geq m(\chi)-2$, and thus that $\operatorname{lk}_{Y}^{h\uparrow}(v)$ is $(m(\chi)-2)$-connected. ∎
@@ 4.2 Exclusion
For any non-zero character $\chi\in\operatorname{Hom}(P\mathcal{H}_{n},\mathbb{R})$, we have $\chi]\notin\Sigma^{m(\chi)}(\mathcal{H}_{n})$.
@@
Let $X$ be a CW-complex covered by subcomplexes $(X_{i})_{i\in I}$ and let $L$ be the nerve of the cover. Let $n\geq 1$. Suppose that any non-empty intersection $X_{i_{1}}\cap\cdots\cap X_{i_{r}}$ is (nr)(n-r)-connected. Then $H_{k}(X)\cong H_{k}(L)$ for all $k\leq n-1$, and $H_{n}(X)$ surjects onto $H_{n}(L)$.
@@
Recall that a subcomplex $Z$ of a CAT(0) cube complex $Y$ is locally combinatorially convex if every link in $Z$ of a vertex $z\in Z^{(0)}$ is a full subcomplex of the link of $z$ in $Y$, and combinatorially convex if it is connected and locally combinatorially convex. It is known that combinatorially convex implies CAT(0), hence contractible. In particular, this applies to connected components of locally combinatorially convex subcomplexes.
@@
It suffices to show that each $X^{\chi_{i}\leq 0}$ is locally combinatorially convex. Given a pair of adjacent vertices, we can write them as $v=[L_{g},\varphi]$ and $w=[L_{g+1},\varphi\circ\rho_{j}^{-1}]$ for some $j$. Then $\chi_{i}(w)-\chi_{i}(v)=\delta_{i,j}$. Thus, if $C$ is a cube containing $v$, and $w_{1},\dots,w_{k}$ are the vertices of $C$ adjacent to $v$, then the maximum and minimum values of $\chi_{i}$ on $C$ lie in $\{\chi_{i}(v),\chi_{i}(w_{1}),\dots,\chi_{i}(w_{k})\}$. Thus, whenever $v\in X^{\chi_{i}\leq 0}$ and all these $w_{j}$’s lie in the link of $v$ in $X^{\chi_{i}\leq 0}$, then the cube $C$ lies in $X^{\chi_{i}\leq 0}$. This implies that the link of $v$ in $X^{\chi_{i}\leq 0}$ is a full subcomplex of the link of $v$ in $X$. ∎
As in [Zar20], we have the immediate corollary that intersections of blankets are blankets for the union of the sets of indices. What follows is essentially identical to Zaremsky’s approach for the Houghton groups, reproduced here with minor changes for the sake of completeness. Recall now the more general statement of the Morse lemma, Lemma 2.11. For notational convenience, we write $X_{f\leq k}$ for $X^{f\leq k}$, and $X_{f\leq k}^{t\leq\chi}$ for the intersection $X_{f\leq k}\cap X^{t\leq\chi}$.
@@
If $X_{f\leq 3n-2}^{0\leq\chi}$ is not $(m(\chi)-1)$-connected, then $\chi]\in\Sigma^{m(\chi)}(P\mathcal{H}_{n})^{c}$.
@@
If $\chi]\in\Sigma^{m(\chi)}(P\mathcal{H}_{n})$, then the filtration $(X_{f\leq 3n-2}^{t\leq\chi})_{t\in\mathbb{R}}$ is essentially $(m(\chi)-1)$-connected. Every $h$-ascending link of a vertex in $X_{f\leq 3n-2}$ is $(m(\chi)-2)$-connected, so for any $s\leq t$ the inclusion $X_{f\leq 3n-2}^{t\leq\chi}\hookrightarrow X_{f\leq 3n-2}^{s\leq\chi}$ induces an isomorphism in $\pi_{k}$ for $k\leq m(\chi)-2$, and a surjection in $\pi_{m(\chi)-1}$. By assumption, for any $t$ there is some $s\leq t$ such that this inclusion induces a trivial map in $\pi_{k}$ for $k\leq m(\chi)-1$, implying that $X_{f\leq 3n-2}^{s\leq\chi}$ is $(m(\chi)-1)$-connected. Rescaling if necessary, we can assume that $s\in\chi(P\mathcal{H}_{n})$, and thus we can translate to obtain $X_{f\leq 3n-2}^{s\leq\chi}\cong X_{f\leq 3n-2}^{0\leq\chi}$, so that $X_{f\leq 3n-2}^{0\leq\chi}$ is itself $(m(\chi)-1)$-connected. ∎
@@
To be non-empty, such an intersection can include at most one term $Y_{i}^{\alpha}$ for each $i$, and can thus be written $Y=Y_{i_{1}}^{\alpha_{1}}\cap\cdots\cap Y_{i_{r}}^{\alpha_{r}}$, with the $i_{j}$ all pairwise distinct. Let $Z=Z_{i_{1}}^{\alpha_{1}}\cap\cdots\cap Z_{i_{r}}^{\alpha_{r}}$, so that $Y=Z\cap X_{f\leq 3n-2}^{0\leq\chi}$. We apply Morse theoretic techniques to $Z$, this time using Lemma 2.11. As $Z$ is an intersection of blankets, it is a blanket, and thus contractible. Given adjacent vertices $w=[L_{g},\varphi]$ and $v=[L_{g+1},\varphi\circ\rho_{i}^{-1}]$ with $v\in Z$, we have that $w\in Z$. Thus, for any vertex of $Z$, the entire $f$-descending link is in $Z$. As this is (n2)(n-2)-connected for $f(v)\geq 2n$, we see that $Z_{f\leq 3n-2}$ is (n2)(n-2)-connected, and so is certainly $(m(\chi)-2)$-connected. Now consider $Y$ as $Z_{f\leq 3n-2}^{0\leq\chi}$. As before, the $h$-ascending link of a vertex $v$ is a join between its $f$-ascending and $f$-descending parts. The latter is in $Z$ for the same reasons as above; the former is in $Z$ because it consists of directions $i$ where $m(\chi)+1\leq i\leq n$, on which the considered $\chi_{i_{j}}$ are constant. Thus, the $h$-ascending link of $v$ is in $Z_{f\leq 3n-2}$. As $Z_{f\leq 3n-2}$ is $(m(\chi)-2)$-connected, the Morse lemma tells us that $Y$ is $(m(\chi)-2)$-connected. ∎
Let $L$ be the nerve of the covering of $X_{f\leq 3n-2}^{0\leq\chi}$ by the $Y_{i}^{\alpha}$. Since $\chi]\in\Sigma^{m(\chi)-1}(P\mathcal{H}_{n})$, we know that $X_{f\leq 3n-2}^{0\leq\chi}$ is $(m(\chi)-2)$-connected, so by the Strong Nerve Lemma $L$ is $(m(\chi)-2)$-connected. The final missing piece is to prove that $L$ is not $(m(\chi)-1)$-acyclic.
@@
It remains to show that $Y_{i}^{1}\neq Y_{i}^{2}$ for each $i$. It suffices to show that $Z_{i}^{1}\neq Z_{i}^{2}$. If these are equal (call them $Z_{i}$), then we can connect $v_{0}$ to $v_{i}$ via a path in $Z_{i}$. In $X$, one could assume that such a path is one on which $f$ first strictly increases, then strictly decreases; since $Z_{i}$ is combinatorially convex, this property holds for $Z_{i}$ as well. Since the path lies in $Z_{i}$, $\chi_{i}$ is non-positive on the whole path. By the uniqueness of ascending edges, this is a path of the form
$Z,\operatorname{id}]$$Z^{\prime},\varphi_{i}]$$\mathcal{O},\operatorname{id}]$$\mathcal{O},\varphi_{i}]$}}